Fermi National Laboratory
 

Operation Summary for August 3, 2011 - August 5, 2011

 

Wednesday Aug. 3

The day shift began with the Tevatron (TEV) off for cryo systems and vacuum systems recovery, with the Recycler off, with the Antiproton Source (Pbar) stacking (33.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments off, in access, and conducting maintenance.

During the day shift, operators reported at 9:59 AM an indication of a Pbar LCW leak. Operators conducted a search of Pbar service buildings, but found nothing. Operators prepared the F manholes for access. Operators began tuning MI at 1:07 PM. Operations put Pbar into standby mode at 3:20 PM.

During the evening shift, operators conducted a MI coalescing check at 4:11 PM. Operators transferred antiprotons from the Accumulator to the Recycler at 4:29 PM and then prepared Pbar for an access to look for a LCW leak. Operators turned on and exercised the TEV Separators at 5:38 PM. Operations established beam to the new Meson FTBF T-1010 experiment at 5:55 PM. Operators reported at 5:59 PM that they found a magnet hose leak in Pbar; Water Works personnel replaced the hose. Operators reported at 7:33 PM that the Pbar horizontal and vertical amplifiers had tripped off; Operators and a Pbar person accessed and replaced a 24-volt power supply for the PLC. Pbar then stacked for 30 minutes and operators put the machine back in standby. Operations disabled beam to T-1010 at 8:07 PM.

Thursday Aug. 4

The midnight shift began with the TEV off for cryo and vacuum systems recovery, with Operations monitoring the stash (80.2E10) and the Pbar stack (9.9E10), and with the NuMI and MiniBooNE experiments taking beam.

During the midnight shift, operators reported at 2:22 AM that the TEV sector C2 air compressor tripped off for the first of three times in about 3.5 hours; it reset. Operators reported at 2:40 AM that the TEV sector F4 air conditioner compressor tripped off for the first of three times in about 2.5 hours; it reset. Operations established beam to T-1010 at 7:14 AM. The Neutron Therapy Facility (NTF) began turning on at 7:21 AM to treat patients. Operators turned off the TEV Separators at 7:34 AM and prepared the TEV for access.

During the day shift, at 9:20 AM, the MI Power Supply interlock Loop (MIPSL) tripped off when the quadrupole VCB at MI-30 opened; MI personnel investigated, could find nothing wrong, and closed the VCB. Operators prepared the TEV transfer hall for vacuum work at 9:23 AM; this access shut off beam to the Meson FTBF experiment T-1010. Operators notified a Pbar person at 9:53 AM that a kicker (PBKICK) tripped off. NTF treated its last patient for the day at 10:43 AM. Pbar began stacking at 1:58 PM. At 2:05 PM, a Booster RF station (BRF7) tripped off for the first of five times in 14 minutes; a Booster RF person investigated. Operations began sending beam to MTA at 2:13 PM. Operators reported that the TEV Transfer Hall came out of access at 3:30 PM.

During the evening shift, Operations established beam to T-1010 at 4:22 PM. Operations disabled beam to MTA at 5:16 PM. Pbar stopped stacking and operators put the machine into standby mode. Operations disabled beam to T-1010 at 10:14 PM.

Friday Aug. 5

The midnight shift began with the TEV off for cryo and vacuum systems recovery, with Operations monitoring the stash (118.3E10) and the Pbar stack (99.6E10), with the NuMI and MiniBooNE experiments taking beam.

During the midnight shift, at 2:13 AM, operators reported a power glitch that took down Linac, Booster, and MI. A Linac quadrupole power supply and a Klystron RF station tripped off. The Booster East Anode and West Anode Power Supplies tripped off. The MI PSL tripped off. A Pbar Debuncher Lambertson tripped off. Three MiniBooNE quadrupole power supplies tripped off. The MINOS near detector tripped off. Operations established beam to NuMI and MiniBooNE at 2:29 AM. FESS personnel reported at 3:03 AM that there were problems with two transformers at the Master Substation; they contacted the High Voltage Coordinator; latter the Coordinator reported that the power glitch was due to ComEd. Operators turned off Booster and MI in preparation for accesses. Operators prepared the TEV Transfer Hall for access at 7:45 AM; again, T-1010 gets shut off due to this access. NTF began preparing at 7:45 AM to treat its only patient for the day.

The Plans for Friday and the Weekend

The plans for today are to continue with recovery for the TEV cryo systems and the vacuum. Continue with Pelletron work. The TEV cryo and vacuum work should be done by noon today. TEV personnel will conduct studies this afternoon and then, hopefully, have their machine ready to establish a store later this evening.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/05/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory